You have given a function λ : R → R with the following properties (x ∈ R, n ∈ N):

λ(n) = 0 , λ(x + 1) = λ(x) , λ(n+1/2)=1
Find two functions p, q : R → R with q(x)6=0 for all x such that λ(x) = q(x)(p(x) + 1).
1
Expert's answer
2020-11-17T17:21:34-0500

The required function is,

"\lambda(x) = \dfrac{1-cos(2πx)}{2}"



Comparing the function to "\lambda(x) = q(x)(p(x) + 1)"


"\therefore p(x) = -cos(2πx),\\
and\ q(x) =\frac{1}{2}"



For x = 1,

"\lambda(1) = \dfrac{1-cos(2π(1))}{2}"


"\lambda(1) = \dfrac{1-cos(2π)}{2} = 0"


For x= 2,

"\lambda(2) = \dfrac{1-cos(2π(2))}{2}"


"\lambda(2) = \dfrac{1-cos(4π)}{2} = 0"



"\therefore" when x = n,

λ(x) = λ(n) = 0


and λ(n) derived is;


"\lambda(n) = \dfrac{1-cos(2π(n))}{2}"


"\lambda(n) = \dfrac{1-cos2nπ}{2}"